OpenStudy (anonymous):

i need to factor 2x^-16

OpenStudy (anonymous):

is this written as 2x^-16 or x^2-16

OpenStudy (anonymous):

@ 8 erica0427

OpenStudy (anonymous):

it is written as 2x^3-16

OpenStudy (anonymous):

@erica0427

OpenStudy (anonymous):

Ok..

OpenStudy (anonymous):

2(x−2)(x2+2x+4)

OpenStudy (anonymous):

\[ 2x^3-16 = 2 [ x^3-8] = 2[x^3-(2)^3] \]Did you understood these steps?

OpenStudy (anonymous):

\[2[x^3-2^3]= 2(x-3)(x^2+2x+2^2) =2(x-3)(x^2+2x+4)\]

OpenStudy (anonymous):

@erica0427

OpenStudy (anonymous):

@dpasingh yes what about this one factor 9z^2-6z-8

OpenStudy (anonymous):

\[9z^2-6z-8 = 9z^2-12z+6z-8= 3z(3z-4)+2(3z-4)= (3z+2)(3z-4)\] are the required factors of the given expression.

OpenStudy (anonymous):

@erica0427

OpenStudy (anonymous):

@erica0427 Did you get this one